OpenStudy (anonymous):

solve the system by substitution -x - y - z = -8 -4x + 4y + 5z = 7 2x + 2z = 4

OpenStudy (anonymous):

so manipulate the first equation to make it x = something

OpenStudy (anonymous):

i have no idea how to get rid of the -y and -z

OpenStudy (anonymous):

Cry

OpenStudy (anonymous):

but its using substitution not elimination

OpenStudy (darkbluechocobo):

I am going to take a gander and say that you are in Algebra 1

OpenStudy (anonymous):

no algebra 2, i think its suppose to be a review but i never learned this.

OpenStudy (darkbluechocobo):

Omy I swear if i have to learn this again o_o

OpenStudy (darkbluechocobo):

But yes this is review from Algebra 1 I hated these things

OpenStudy (anonymous):

im doing online school this year so they probably have it different i dont kknow hahah

OpenStudy (darkbluechocobo):

Haha I am too but I have been doing it for like 2 years now

OpenStudy (anonymous):

if you rearrange the first formula then you can use it to substitute into the next ones

OpenStudy (anonymous):

yeah im failing haha. i hate it.

OpenStudy (darkbluechocobo):

Omy o_o But I must say I am sorry I do not know substitution ._.

OpenStudy (darkbluechocobo):

Sad face

OpenStudy (anonymous):

its alright. neither do i ;)

OpenStudy (darkbluechocobo):

Ding Ding Ding I remember

OpenStudy (darkbluechocobo):

So you are given 3 equations

OpenStudy (anonymous):

yes

OpenStudy (darkbluechocobo):

basically you have to solve one of them and we will go from there

OpenStudy (darkbluechocobo):

the question is which ._. because your school was so nice and give us 3 variables... yay

OpenStudy (anonymous):

i know i cant find any examples that have 3 :/

OpenStudy (darkbluechocobo):

i did :p so one moment

OpenStudy (darkbluechocobo):

Found out how :3

OpenStudy (anonymous):

how??

OpenStudy (darkbluechocobo):

sigh but none of yours cancels out ┻━┻ ︵ヽ(`Д´)ノ︵ ┻━┻

OpenStudy (darkbluechocobo):

idea take the 1st to equations and multiply them by eachother

OpenStudy (anonymous):

ill just be like this question is stupid lol

OpenStudy (anonymous):

see what grade i get hahaha

OpenStudy (darkbluechocobo):

so -5x+3y+4z=-1

OpenStudy (anonymous):

o-O how

OpenStudy (darkbluechocobo):

because you added the 1st line and the second line -x - y - z = -8 -4x + 4y + 5z = 7

OpenStudy (anonymous):

how is substitution different than elimination

OpenStudy (darkbluechocobo):

because basically you are going to have to substitute it into a equation for x y or z depending on the situation

OpenStudy (darkbluechocobo):

But sadly my mind is tangled I have never had to deal with more than 2 variables

OpenStudy (darkbluechocobo):

here is a video that may help but I don't know http://www.youtube.com/watch?v=a3S2NQJYI5o

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

I can help if you still have problems

OpenStudy (darkbluechocobo):

Yes please do because this is tangling my brain ._. I can only do 2 variables not 3

OpenStudy (anonymous):

ok so start with re arranging the fist formula so you have x = what do you get

OpenStudy (anonymous):

are you having problems?